### v2

# Supporting the development of Machine Learning for fundamental science in a federated Cloud with the AI_INFN platform

Abstract
Machine Learning (ML) is driving a revolution in the way scientists design, develop, and deploy data-intensive software. However, the adoption of ML presents new challenges for the computing infrastructure, particularly in terms of provisioning and orchestrating access to hardware accelerators for development, testing, and production. The INFN-funded project $\mathrm{AI\_INFN}$ (“Artificial Intelligence at INFN”) aims at fostering the adoption of ML techniques within INFN use cases by providing support on multiple aspects, including the provision of AI-tailored computing resources. It leverages cloud-native solutions in the context of INFN Cloud, to share hardware accelerators as effectively as possible, ensuring the diversity of the Institute’s research activities is not compromised. In this contribution, we provide an update on the commissioning of a $\mathrm{Kubernetes}$ platform designed to ease the development of GPU-powered data analysis workflows and their scalability on heterogeneous, distributed computing resources, possibly federated as Virtual Kubelets with the interLink provider.
1 Introduction
Over the past decade, Artificial Intelligence (AI) has seen rapid and widespread adoption, establishing itself as a standard tool for processing large, complex datasets, and extracting insights from multi-modal, multi-domain data Bommasani2021FoundationModels . The proliferation of text-to-image apps and the advent of AI-powered chat-bots have helped propel AI into the mainstream, consolidating its explosion in both usage and development. Today, AI is reshaping the computing landscape, driving technological evolution, influencing hardware market trends, and dominating software development worldwide.
Cloud computing has also played a key role in accelerating the adoption of AI techniques by making computing power and data storage more accessible to developers by outsourcing resource management to service providers like AWS AWS or GCP GCP . This approach enables ready-to-use software to be delivered directly to the end users, a model known as Software-as-a-Service (SaaS). A typical example is JupyterHub JupyterHub , which provides multiple users with access to notebooks for data visualization and interactive computation using resources provisioned by the Cloud. The growing community of Data Scientists makes large use of this or similar technologies, contributing to expanding the catalog of libraries and tools available for developing and deploying AI applications.
The scientific community is closely following the evolution of Machine Learning (ML), aiming at adapting advanced ML-based algorithms for fundamental research. This is particularly true in the High Energy Physics (HEP) field, where researchers are exploring AI-driven computing solutions to accelerate the workflow, from data acquisition and simulation to physics analysis Albertsson:2018maf . The paradigm of interactive computing is also gaining interest within the HEP community as a promising solution to meet the growing resource demand expected from next-generation experiments. These efforts are leading to the design and implementation of Analysis Facilities (AFs), a collection of infrastructures and services that integrates data, software, and computational resources to execute one or more elements of an analysis workflow Ciangottini:2024vtl .
The Italian National Institute for Nuclear Physics (INFN) is the coordinating institution for nuclear, particle, astroparticle, medical and theoretical physics in Italy. It promotes, coordinates, and conducts scientific research, along with the technological developments needed for the activities in these fields. In response to the paradigm shift driven by AI and Cloud technologies, INFN is reorganizing its computing infrastructure to support emerging trends, strengthening its resources Grandi:2024fqt , and expanding the range of services offered through INFN Cloud [^1] Cloud resources for INFN research, https://www.cloud.infn.it. Salomoni:2024dft . Within this context, the $\mathrm{AI\_INFN}$ (“Artificial Intelligence at INFN”) initiative was launched, aiming at sharing hardware resources, learning best practices, and developing AI-powered applications relevant to INFN research. In this document we present the $\mathrm{AI\_INFN}$ platform and discuss its relevance to the outlined scopes.
2 The AI_INFN initiative
The $\mathrm{AI\_INFN}$ initiative, aims to connect the scientific communities developing infrastructures, algorithms, and applications. It provides tools, computing infrastructures, and social connections to foster collaboration and facilitate the adoption of AI-powered computing technologies within INFN research fields.
$\mathrm{AI\_INFN}$ is organized in four work packages: procuring and operating computing infrastructure with hardware accelerators, organizing training events for Machine Learning adoption, building a community of ML experts and developers across INFN units, and developing the competence to profit from hardware accelerators beyond Graphics Processing Units (GPU), such as FPGAs and Quantum Processors, for AI model training and inference.
[…]
The $\mathrm{AI\_INFN}$ platform is a SaaS provided by $\mathrm{INFN\ Cloud}$. The underlying specialized and dedicated hardware was designed for High Performance Computing (HPC) tasks and is hosted and managed at INFN CNAF in Bologna, few steps away from the pre-exascale CINECA supercomputer Leonardo [^2] More details available at https://leonardo-supercomputer.cineca.eu. The infrastructure comprises four servers, clustered in an OpenStack OpenStack tenancy, acquired and installed between 2020 and 2024 as the demand for Cloud-accessible HPC computing resources increased:
  • Server 1 (2020), with 64 CPU cores, 750 GB of memory, 12 TB of NVMe disk, eight NVIDIA Tesla T4 GPUs and five NVIDIA RTX 5000 GPUs;
  • Server 2 (2021), with 128 CPU cores, 1024 GB of memory, 12 TB of NVMe disk, two NVIDIA Ampere A100 GPU, one NVIDIA Ampere A30 GPU, two AMD-Xilinx U50 boards and an AMD-Xilinx U250 board;
  • Server 3 (2023), with 128 CPU cores, 1024 GB of memory, 24 TB of NVMe disk, three NVIDIA Ampere A100 GPUs and five AMD-Xilinx U250 boards;
  • Server 4 (2024), with 128 CPU cores, 1024 GB of memory, 12 TB of NVMe disk, one NVIDIA RTX 5000 GPUs and two AMD-Xilinx Versal V70 boards.
Before $\mathrm{AI\_INFN}$ started its activity in January 2024, the farm was maintained by another INFN initiative named $\mathrm{ML\_INFN}$ that, created as a proof-of-concept for designing a platform for sharing accelerated resources, was developed with a provisioning model relying on Virtual Machines (VMs) assigned to groups of users developing a data analysis or Machine Learning study Anderlini:2024oup . During the late period of $\mathrm{ML\_INFN}$, however, an increase in the user base highlighted some limitations to the efficiency of this provisioning model. These limitations were related to administrative and user-support burden, very long idling times, and dangerous eviction of the stateful user’s deployments. In 2023, the security risks grew to an unacceptable level and called for the introduction of an alternative model enabling users to tune the resources provisioned to their cloud-based computing environment, without becoming administrators of a multi-user web service.
At the time of writing, 72 researchers working on 16 research activities have requested and gained access to the platform. On average, 10 to 15 researchers connect at least once to the platform in a working day. Two dedicated clones of the $\mathrm{AI\_INFN}$ platform were temporarily deployed at CNAF and at ReCaS Bari to provision the GPU-accelerated resources to the 30 participants of the first AI_INFN hackathon, an advanced training event organized in Padua in November 2024333More details available at https://agenda.infn.it/event/43129..
3 SaaS provisioning model, the AI_INFN platform
The $\mathrm{AI\_INFN}$ platform was deployed on a $\mathrm{Kubernetes}$ Kubernetes cluster spanning on at least three VMs within the dedicated OpenStack tenancy providing part of the storage resource, the monitoring infrastructure and the $\mathrm{Kubernetes}$ control plane. A minimal amount of compute resources is also provisioned to make it possible for users to access their data on the platform anytime Anderlini:wmlq2024 . Additional compute resource provided by VMs can be attached to the cluster and detached to be used as standalone machines running an Ansible Ansible playbook, or reassigned to another cluster in the same tenancy. $\mathrm{AI\_INFN}$ users are identified through INFN Cloud Indigo IAM indigoiam instance. Once authenticated, users can configure and spawn their $\mathrm{JupyterLab}$ instance using $\mathrm{JupyterHub}$.
The main platform file system is distributed through the containers via NFS. One of the platform nodes runs an NFS server in a $\mathrm{Kubernetes}$ pod and exports data to the containers spawned by $\mathrm{JupyterHub}$. At spawn time, $\mathrm{JupyterHub}$ is configured to create the user’s home directories and project-dedicated shared volumes. A special directory of the platform file system, that users can use directly or clone and extend in their directories, is reserved for distributing managed software environments, configure using virtual environments. The platform file system is subject to regular encrypted backup. Backup data is stored in a remote Ceph ceph volume provisioned by $\mathrm{INFN\ Cloud}$ using the BorgBackup BorgBackup package to ensure data deduplication.
[…]
To address the bandwidth limitations of a virtual file system with a remote backend, which can hinder iterative training and data analysis requiring to process the whole dataset multiple times, the $\mathrm{AI\_INFN}$ platform provides also an ephemeral file system. This system is mapped directly to a logical volume on the hypervisor’s NVMe storage. The indications for the users is to copy the required data to this fast volume at the beginning of each session. These ephemeral volumes are also useful as a cache for intermediate results or to extend RAM through memory mapping.
At the opposite extreme of the I/O performance spectrum there are distributed virtual file systems that can be mounted on multiple computing resources, enabling the sharing of notebooks and user-defined computing environments across multiple computing sites and compute backends. JuiceFS JuiceFS is a cloud-based, high-performance, POSIX-compliant distributed file system specifically designed for multi-cloud and serverless computing. It decouples data and metadata delegating these tasks to highly optimized third party projects, combining a metadata engine implemented with either key-value databases (such as Redis redis-link ) or relational database management systems (such as PostgreSQL PostgreSQL ) with storage systems accessed through S3, WebDAV or other high-throughput protocols.
Finally users can install or upgrade packages in their containers. Installing new software will introduce ephemeral modifications in OverlayFS layer on top of the container file system. A more effective and popular alternative to installing packages in the container is to rely on the binaries distributed through the CERN VM file system (cvmfs) Buncic:2010zz . CVMFS, used to distribute software through the nodes of the WLCG, is made available to the platform users through a $\mathrm{Kubernetes}$ installation that shares the caches among different users and sessions.
In the $\mathrm{AI\_INFN}$ platform, the NVIDIA GPU operator GpuOperator , is used to install and maintain the GPU drivers. The NVIDIA GPU Operator on $\mathrm{Kubernetes}$ streamlines the management and deployment of NVIDIA GPU resources by automating the installation and configuration of required components within a $\mathrm{Kubernetes}$ cluster. In general, the GPU Operator is designed to simplify the management of clusters with a large number of GPU accelerators, enabling efficient scaling and resource optimization for GPU-accelerated workloads such as AI, ML, and data processing. This approach enables centralized management of GPUs, ensuring a consistent and scalable configuration across all nodes in the cluster, while simplifying maintenance and updates.
One of the most common support requests with the VM-based provisioning model involved setting up a GPU-accelerated Python software stack. While the TOSCA template TOSCA and Ansible playbook handled the installation of the NVIDIA driver and runtime, choosing and installing the Python libraries required for the application was left to the users.Typically, managing a data science project’s dependencies is the responsibility of developers and analysts, and many analysis projects require multiple computing environments. When introducing the $\mathrm{AI\_INFN}$ platform, particular attention was given to ensuring user sessions were highly customizable and adaptable by providing mechanisms for users to create and manage their own computing environments. The most radical customization option is to build and pick a custom OCI image. Both communities and individual users can modify the default OCI image by adding system libraries or software packages or by altering the $\mathrm{JupyterLab}$ service itself.While users often prefer conda conda for custom software environments, Apptainer Apptainer images are gaining popularity. Unlike conda, which consists of thousands of small files, Apptainer uses SquashFS Squashfs , a compressed read-only file system, to package the entire environment into a single file. This makes Apptainer images easier to share and distribute through object stores.The $\mathrm{AI\_INFN}$ platform provides documentation to help users export conda environments as Apptainer images and use them as Jupyter kernels. It also offers pre-built conda environments and Apptainer images with software versions optimized for GPU-accelerated Machine Learning frameworks. Users can clone these environments and add project-specific dependencies, typically related to data loading and visualization, and independent of the GPU software stack.A notable exception is represented by the software environment to develop Quantum Machine Learning (QML), featuring Python modules that simulate the effect of quantum operators on GPU and therefore requiring the same attention as other GPU-accelerated ML libraries to match the versions of the underlying software.In addition, Apptainer images specialized for the data processing of the LHC experiments can be obtained via CVMFS.
A dedicated monitoring and accounting system has been set up for the platform in order to effectively control the use of all the platform’s resources and in particular of the GPUs. Several metric exporters have been configured to collect the information of interest and then expose it to a Prometheus Prometheus-link instance running in the platform. Some of these exporters are already available as Free and Open Source Software, such as Kube Eagle KubeEagle , which manages information about the use of the cluster’s CPUs and memory resources by the various components of $\mathrm{Kubernetes}$, or NVIDIA GPU DCGM exporter DCGM . Other exporters were developed on purpose, for example to monitor the usage of storage resources. All the metrics collected by Prometheus are then made visible and accessible through a Grafana grafana-docs dashboard. Grafana is run in a VM independent of the platform cluster and is used to monitor other VMs in the $\mathrm{AI\_INFN}$ OpenStack tenancy. It also hosts a PostgreSQL database for the accounting metrics, updated at regular intervals by averaging the metrics obtained from the monitoring Prometheus service.
4 Offloading: scale the applications beyond cluster boundaries
While the $\mathrm{AI\_INFN}$ platform is primarily conceived as the to-go solution for researchers moving their first steps with hardware-accelerated and machine learning software development, it was designed to enable packaging and scaling the developed applications with computing resources made available in remote computing centers with a mechanism known as offloading.
The architecture of the offloading capabilities of the $\mathrm{AI\_INFN}$ platform consists of few self-consistent components that can be used by the application to scale beyond the single notebook instance. Indeed, different components may introduce different limits or limitations on the scaling capabilities and should be selected wisely on a per-application basis.
Users are allowed to scale beyond their notebook instance by creating $\mathrm{Kubernetes}$ jobs, enqueued and assigned to either local or remote resources by the Kueue controller Kueue . Kueue is designed to use local resources in an opportunistic way, configuring the running batch jobs to be immediately evicted in case new notebook instances are spawned pushing the cluster in a condition of resource contention. User do not create jobs directly accessing $\mathrm{Kubernetes}$ APIs, but passing through a dedicated microservice, named vkd, that validates user’s request based on memberships criteria and manage $\mathrm{Kubernetes}$ secrets that are not intended to be exposed to users, but still are needed for their jobs to be executed in the platform.
An interesting feature supported by vkd is the ability of cloning the notebook instance, replacing the start-up commands spawning the notebook with user-defined commands. In practice, these Bunshin Jobs provide a very simple interface to scale the application within the cluster boundaries as the applications developed within the notebook instance are guaranteed to run identically in the cloned instances.
Figure 1: The $\mathrm{AI\_INFN}$ platform enables offloading using InterLink as a virtual Kubelet provider to offload workloads from a Kubernetes cluster to external resources. InterLink interfaces with HTCondor, Slurm, and Podman via dedicated plugins, providing transparent access to HPC and cloud infrastructures.
Figure 2: Scalability test involving: the INFN-Tier1 at CNAF in Bologna provisioned via HTCondor (labeled infncnaf); the CINECA Leonardo super computer in Bologna provisioned via Slurm (leonardo); a Virtual Machine in the Cloud provisioned via Podman (podman); a Terabit HPC-Bubble in Padova provisioned via Slurm (terabitpadova). The label recas in the legend refers to a WLCG Tier-2 site in Bari, integrated, but not taking part to the test.
Users can flag their jobs at submission time as compatible with offloading. The compatibility of a job with offloading should be evaluated considering technical aspects (for example, an offloaded job cannot rely on the local storage resources such as NFS), practical consideration (for example, the longer delay between submission and execution in large data centers may make offloading ineffective for very short jobs), and policy restrictions (for example, secrets to access confidential data cannot be shared with a remote data center). Kueue may then assign jobs marked as compatible with offloading to virtual nodes.
Virtual nodes are $\mathrm{Kubernetes}$ nodes that are not backed by a Linux kernel but mimic a $\mathrm{Kubernetes}$ kubelet in the interactions with the $\mathrm{Kubernetes}$ API server. The software component providing this interface is named Virtual Kubelet vk and is designed to ease the integration with various resource providers. The $\mathrm{AI\_INFN}$ platform relies on the InterLink Ciangottini:chep2024 provider. A further abstraction layer defining a simplified set of REST APIs that can be implemented by the so-called InterLink plugins providing the actual access to the compute resources. At the time of writing, the $\mathrm{AI\_INFN}$ platform is interfaced with plugins accessing HTCondor htcondor , Slurm slurm and Podman podman resources. Following a recent integration test, a $\mathrm{Kubernetes}$ plugin will be brought to production soon. A schematic representation of the architecture is provided in Figure 2.
To ease the deployment of application in the remote data centers, the $\mathrm{AI\_INFN}$ platform relies on dedicated and distributed file system based on JuiceFS using Redis as metadata engine and an S3 endpoint for data storage. The secrets to mount the shared file system are shipped to the remote data center that, if allowed by site-specific policies, can make it available to the applications as a FUSE file system. Relying on the distributed file system drastically hinder the scalability of the developed application, but provides a precious intermediate level between cluster-local development and multi-site distributed production.
Figure 2 reports a recent scalability test involving resources provisioned by four different sites, without distributing the file system and for CPU-only payloads of the LHCb Flash Simulation Barbetti:2906203 . The plot illustrates the increase in job counts and their distribution across multiple computing sites. Stable contributions were maintained throughout the test by the INFN-Tier1 at CNAF in Bologna, provisioned via HTCondor. Comparable stability was observed from a Podman-based VM, whereas CINECAâs Leonardo supercomputer, provisioned via Slurm, showed greater variability in job activity during the initial hours. Later in the test, the Terabit HPC-Bubble, also provisioned via Slurm, was integrated and demonstrated consistent, sustained usage. The pronounced increase in job activity during the evening suggests improved scheduling efficiency. Overall, the test confirms the effectiveness of integrating multiple sites with heterogeneous provisioning mechanisms.
5 Conclusion
Machine Learning and Artificial Intelligence have been reshaping the landscape of data processing and data analysis applications, making it easier for data analysts and data scientist to accelerate a variety of computing-intensive tasks on GPUs. The $\mathrm{AI\_INFN}$ initiative is developing and serving a highly customizable development platform, integrated in the service portfolio of $\mathrm{INFN\ Cloud}$ and provisioning different GPU models, possibly installed in remote computing centers through offloading techniques. Although primarily a research and development project, the $\mathrm{AI\_INFN}$ platform is gaining recognition from several small experiments within the $\mathrm{AI\_INFN}$ research lines as a potential provider of GPU-accelerated computing resources, and is poised to play a trailblazing role in the future landscape of the INFN computing infrastructure.
Acknowledgements
The work presented in this paper is performed in the framework of Spoke 0 and Spoke 2 of the ICSC project – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by the NextGenerationEU European initiative through the Italian Ministry of University and Research, PNRR Mission 4, Component 2: Investment 1.4, Project code CN00000013 - CUP I53C21000340006.